OpenStudy (anonymous):

At 10:30 a.m. two planes leave Houston, one flying east at 560 km/h and the other flying west at 640 km/h. At what time will they be 2100 km apart?

OpenStudy (anonymous):

Since the planes are going in opposite direction and you want a specific distance between the two, you can add both the velocities together then divide by 2100km for the time. Then you will have the time in hours and you can do the math to find what time it will be from 10:30 am. I used the equation V=d/t for this. (V1 + V2) = 2100/t and solve for t.